What is the answer to this, one third times the square root of 63?
Found 2 solutions by stanbon, checkley77:
Answer by stanbon(75887)

What is the answer to this, one third times the square root of 63?
-----------------
(1/3)sqrt(63)
---
= (1/3)sqrt(9*7)
= (1/3)*3*sqrt(7)
= sqrt(7)
================
Cheers,
Stan H.

Answer by checkley77(12844)

1/3(sqrt63)
1/3*7.937
7.937/3
2.646 ans